OpenStudy (khally92):

Help please A solid conducting disc experiences a changing magnetic field of the form B = 0.478*t T, where t is the time measured in seconds and the field is perpendicular to the disc. If on this surface, one were to examine a circular current path of radius 0.056 m, what would be the magnitude of the induced electric field around this path?

OpenStudy (osprey):

I've seen this geometry somewhere before ... this is a "think about it" job and see what I can come up with. Dare I say it ... "wish me luck" ?

OpenStudy (khally92):

....... haha Ok, best of luck . And thank's

OpenStudy (osprey):

If you can get access to a book by Lorrain and Corson on Elec and Mag I THINK that that's where I may have seen it. I think it's got a name, but you've caught me flat footed (just for a change).

OpenStudy (khally92):

Ok I will go through the book, I just downloaded it

OpenStudy (khally92):

Nah I dont think so.

OpenStudy (festinger):

I'll assume you know what magnetic flux and Faraday's law is. From Faraday's law, we have: \[\mathcal{E}=-\frac{\mathrm{d}\Phi_{B}}{\mathrm{d}t} \]where \(\mathcal{E}\) is the induced EMF and \(\Phi_{B}\) is the magnetic flux. Thus, to compute the induced EMF simple evaluate: \[\mathcal{E}=-\pi r^{2}\frac{\mathrm{d}B}{\mathrm{d}t} \]

OpenStudy (khally92):